OpenStudy (anonymous):

Choose the value of the x determinant (Dx) in the following system. 2x – 3y = 5 4x + 2y = 7

OpenStudy (texaschic101):

are you just looking for the value of x ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

|2 -3| |4 2| 2*2 - 4*-3 = 4 - (-12) = 16 ?? Yu might need a second opinion on this....

OpenStudy (texaschic101):

oops...typo...*you

jimthompson5910 (jim_thompson5910):

texaschic101, you found the determinant D but what they want is Dx so you replace the first column with 5 7 then you compute the determinant

OpenStudy (raden):